# Nexus: Higher-Order Attention Mechanisms in Transformers

### v1

@@ Abstract
Transformers have achieved significant success across various domains, relying on self-attention to capture dependencies. However, the standard first-order attention mechanism is often limited by a low-rank bottleneck, struggling to capture intricate, multi-hop relationships within a single layer. In this paper, we propose the Higher-Order Attention Network (Hon), a novel architecture designed to enhance representational power through a recursive framework. Unlike standard approaches that use static linear projections for Queries and Keys, Hon dynamically refines these representations via nested self-attention mechanisms. Specifically, the Query and Key vectors are themselves outputs of inner attention loops, allowing tokens to aggregate global context and model high-order correlations prior to the final attention computation. We enforce a parameter-efficient weight-sharing strategy across recursive steps, ensuring that this enhanced expressivity incurs $\mathcal{O}(1)$ additional parameters. We provide theoretical analysis demonstrating that our method breaks the linear bottleneck of standard attention. Empirically, Hon outperforms standard Transformers on multiple benchmarks.
@@ 1 Introduction
To bridge this gap, we propose a novel Higher-Order Attention Mechanism (Hon) specifically designed to enhance the reasoning capabilities of Transformers. Instead of simple linear projections, our approach recursively refines Query and Key vectors through nested attention loops. This recursive formulation allows the model to capture high-order interactions—effectively performing a ”pre-reasoning” step to align semantic relationships before the final attention computation. While recursive computation inherently increases inference FLOPs, we introduce a Weight-Shared strategy that maintains the same parameter count as standard Transformers. This design presents a strategic trade-off: investing more compute per token to achieve significantly higher reasoning density and expressivity. We demonstrate both theoretically and empirically that this mechanism breaks the low-rank bottleneck. Furthermore, we show that Hon is not just a pre-training architecture but a powerful ”upgrade kit” for existing LLMs, significantly boosting the mathematical reasoning abilities of models like Qwen2.5 through architectural retrofitting.
Figure 1: Overview of the Higher-Order Attention Network (Hon). The figure illustrates the hierarchical structure of our proposed mechanism. Left: The integration of the Hon layer within a standard Transformer block, replacing the conventional self-attention module. Middle: The detailed architecture of the 2nd-Order Attention mechanism. Unlike standard attention where $Q$ and $K$ are linear projections, Hon recursively refines $Q$ and $K$ through inner self-attention loops (MatMul $\rightarrow$ SoftMax $\rightarrow$ MatMul) before the final attention computation. This allows the model to capture intricate dependencies prior to the main interaction. Right: The generalized Recursive Attention framework, demonstrating how the mechanism can be extended to arbitrary orders ($m$-th order) to model deeper hierarchical relationships.
@@ 2.3 Reasoning and Chain-of-Thought in LLMs
The ability of LLMs to perform complex reasoning, often elicited via Chain-of-Thought (CoT) prompting (wei2022chain), has become a focal point of research. While scaling laws (kaplan2020scaling) suggest that larger models generally reason better, architectural limitations remain. Standard self-attention treats tokens as a flat sequence, which may struggle to maintain the rigid logical consistency required for multi-step mathematical or symbolic reasoning (dziri2023faith). Recent works have attempted to bolster reasoning via external verifiers (cobbe2021training) or iterative prompting strategies. In contrast, our work seeks to internalize this iterative reasoning capability directly into the attention mechanism. By recursively refining Queries and Keys, the Hon network effectively performs a coherent ”pre-reasoning” step within the layer itself, potentially reducing the reliance on massive depth for logical deduction.
@@ 3.4 Parameter-Efficient Weight Sharing
where the inner attention uses the same projection weights as the outer loop. This constraint ensures that the parameter complexity of our Hon network remains identical to that of a standard Transformer, i.e., $\mathcal{O}(1)$ with respect to the recursive order $m$. As demonstrated in our ablation study, this strategy maintains high performance while maximizing parameter efficiency.
@@ 3.5 Complexity Analysis
Crucially, thanks to our Weight-Sharing strategy, the parameter complexity remains $\mathcal{O}(1)$ with respect to $m$. This makes Hon highly memory-efficient during training compared to simply stacking more Transformer layers, as it does not increase the model size (storage), only the computational density.
@@ 3.6 Linear Bottleneck of Standard Attention Mechanism
Table 1: Main Results: Zero-shot accuracy comparison between Pythia baselines and our proposed Hon networks across different model scales. The best performance for each scale is highlighted in bold.
@@
As depicted in Figure 2, the attention matrices of the Hon network exhibit more intricate and interconnected patterns compared to the standard Transformer and other higher-order transformers. These richer patterns indicate that the Hon network effectively captures multifaceted relationships among tokens, facilitating enhanced reasoning and contextual understanding. Specifically, the Hon network’s attention matrices demonstrate a higher degree of connectivity and diversity in attention weights, reflecting its ability to model complex, hierarchical dependencies more comprehensively.
@@ 4.1 Evaluation on Different Scales
We adhered to the exact training settings employed by Pythia, including learning rates, optimizers, and other hyperparameters, and utilized the Pile dataset. The Pile (gao2020pile) is an 825 GiB corpus of English text, specifically designed for training large-scale language models. This project is composed of 22 distinct, high-quality subsets, both pre-existing and newly constructed, many of which originate from academic or professional sources. This comprehensive and diverse dataset serves as a robust foundation for developing and fine-tuning language models Our Hon model was trained with the same setting as pre-trained Pythia model. We evaluated our approach on six public datasets used by Pythia: PIQA (bisk2020piqa), Hellaswag (zellers2019hellaswag), Sciq (welbl2017crowdsourcing), ARC-E, ARC-C (clark2018think), and LogiQA (liu2020logiqa).
Table 1 summarizes the results. The Hon network outperforms the standard Transformer baseline in terms of average accuracy across all model scales. Notably, our method achieves substantial gains on tasks requiring multi-step reasoning or long-context integration, such as SciQ (+6% on 70M) and PiQA. While the baseline Pythia model occasionally performs slightly better on retrieval-heavy tasks like ARC-C at specific scales (e.g., 1B), Hon demonstrates a more robust capability in handling complex logical dependencies, leading to superior overall performance. This aligns with our hypothesis that higher-order attention is particularly beneficial for tasks where the relationship between tokens cannot be captured by simple pairwise interactions.
@@ 4.2 Ablation Study
To thoroughly validate the design choices of our Higher-Order Attention Network (Hon), we conducted a series of ablation studies on the 70M parameter scale. We utilized the same training configuration as the Pythia-70M baseline. Our analysis focuses on three key aspects: (1) the selection of projections for higher-order attention, (2) the impact of parameter sharing on efficiency, and (3) the effect of recursive order depth. The results are summarized in Table 2.
@@
Hon-QK-Shared Q, K Yes 2 0.206 0.357 0.367 0.273 0.601 0.630 0.406
Hon-Recursive Q, K Yes 3 0.201 0.391 0.368 0.287 0.582 0.659 0.415
@@
We first investigated which components of the attention mechanism ($Q$, $K$, $V$) benefit most from higher-order processing. As shown in Table 2, applying higher-order attention only to the Query projection (Hon-Q) yields a marginal improvement over the baseline (0.400 vs. 0.397). However, extending the mechanism to both Query and Key (Hon-QK) results in a significant performance boost, raising the average accuracy to 0.409. Interestingly, further applying the mechanism to the Value projection (Hon-QKV) does not provide additional gains compared to Hon-QK (both achieve $\sim$0.409). This suggests that the core benefit of our method lies in refining the correlation alignment between Queries and Keys, while a simple linear projection suffices for Values. Consequently, we adopted the Q and K configuration for subsequent experiments.
@@
While Hon-QK achieves superior performance, the introduction of separate attention layers for projections increases the parameter count. To improve parameter efficiency, we proposed a weight-sharing strategy (Hon-QK-Shared), where the parameters of the inner attention layers are shared with the outer main attention layer. Comparing Hon-QK and Hon-QK-Shared, we observe a slight performance drop (0.409 $\rightarrow$ 0.406), yet the shared model still consistently outperforms the baseline. This trade-off allows us to maintain the high expressiveness of higher-order attention while keeping the parameter count comparable to the standard Transformer, maximizing parameter utilization.
@@
Finally, we explored the potential of increasing the recursive depth. By extending the shared model to a 3rd-order attention (Hon-Recursive), the performance further improves to 0.415, surpassing all other configurations. This result empirically validates our theoretical claim that higher recursive orders can capture more intricate dependencies. However, increasing the order also incurs higher computational costs during inference. Considering the balance between efficiency and accuracy, we selected the 2nd-order shared configuration (Hon-QK-Shared) as the default setting for our main experiments, as it offers a robust improvement over the baseline with optimal resource efficiency.
@@ 4.3 Visualization of Attention Patterns
To intuitively understand how the Higher-Order Attention mechanism processes information differently from the standard Transformer, we visualize the average attention weights across all layers and heads for the 70M scale models. We compare the baseline Pythia model with our proposed Hon model. For the Hon model, we examine three distinct attention maps: the distinct inner recursive attentions for Queries ($\text{Hon}_{Q}$) and Keys ($\text{Hon}_{K}$), and the final outer attention ($\text{Hon}_{\text{outer}}$). The visualization results are presented in Figure 2.
@@
(b) Hon Outer
(c) Hon Inner ($Q$-Attn)
(d) Hon Inner ($K$-Attn)
Figure 2: Visualization of average attention heatmaps. The x-axis represents Key positions, and the y-axis represents Query positions. Brighter colors indicate higher attention weights. (a) Standard Self-Attention in Pythia. (b) The outer main attention of the Hon network. (c) The inner recursive attention used to project Queries. (d) The inner recursive attention used to project Keys.
@@
Comparing the baseline attention (Figure 2a) with the Hon outer attention (Figure 2b), we observe that both maintain the fundamental structural characteristics of causal language models: a strong diagonal focus (attending to local context) and a prominent vertical band on the left (attending to the beginning-of-sentence token). This similarity is crucial; it indicates that the Higher-Order Attention mechanism preserves the essential capabilities of the Transformer to model sequential order and global context, ensuring stability while enhancing expressiveness.
@@
In a standard Transformer, the Query and Key projections are linear and context-agnostic (i.e., $K_{i}$ depends only on $x_{i}$). In contrast, our visualization confirms that in the Hon network, the vector used as a ”Key” for the outer loop is itself a context-aware representation aggregated from previous tokens via the inner loop. The visualization of $\text{Hon}_{Q}$ and $\text{Hon}_{K}$ demonstrates that the model learns to dynamically adjust these projections based on the sequence context, effectively performing a ”pre-reasoning” step that simplifies the task for the final outer attention layer.
@@ 4.4 Retrofitting Standard Transformers for Reasoning
To demonstrate the versatility and practical applicability of our approach, we investigated whether existing standard Transformer models could be ”upgraded” to the Hon architecture during the Supervised Fine-Tuning (SFT) stage. We utilized the Qwen2.5-Base models (1.5B and 7B) as our starting points.
@@
Table 3: Performance comparison of Qwen2.5 models fine-tuned with standard architecture vs. retrofitted Hon architecture on reasoning benchmarks. SFT denotes standard Supervised Fine-Tuning, while Hon-SFT denotes fine-tuning after converting the architecture. Scores are reported as accuracy.
@@
  • On the 1.5B scale, Hon-SFT achieves a significant gain on MATH-500 (+1.5%) and improvements on GPQA, demonstrating that smaller models can benefit from the enhanced expressiveness of higher-order interactions to solve math problems.
  • On the 7B scale, the improvements are particularly notable in the AIME24 benchmark (+2.3%), which consists of challenging mathematics competitions. This suggests that as the model scale increases, the Hon architecture effectively leverages the recursive mechanism to handle complex, multi-step reasoning chains required for competition-level mathematics.
These findings suggest that the Hon architecture is not limited to pre-training from scratch but can serve as an effective ”architectural upgrade” for existing models, unlocking further potential in complex reasoning tasks with minimal adaptation cost.
@@ 5 Conclusion
Transformers have achieved significant success across various domains, primarily due to their self-attention mechanisms that efficiently capture long-range dependencies. However, traditional first-order attention methods encounter limitations when addressing complex tasks involving intricate, multi-token relationships. In this paper, we introduce the Higher-Order Attention Network (Hon network), which employs a novel recursive higher-order attention mechanism to process query and key vectors iteratively, thereby constructing a multi-level attention framework. This approach enhances the model’s ability to capture hierarchical dependencies and complex interactions within a single attention layer, leading to improved scalability and performance. Theoretically, we demonstrate that our higher-order mechanism offers greater expressiveness compared to existing methods. Empirical evaluations on multiple benchmark datasets confirm that the Hon network consistently outperforms standard Transformers. Moreover, we demonstrated that our architecture allows for the effective upcycling of existing pre-trained models. By retrofitting standard Transformers with Higher-Order Attention during the fine-tuning stage, we achieved significant gains in complex mathematical reasoning benchmarks. These advancements position the Hon network not only as a robust foundation for new models but also as a practical enhancement for the current generation of Large Language Models. Future work will focus on optimizing the computational efficiency of higher-order attentions and exploring their applicability in broader domains, including vision and multimodal learning.

### v2

@@ Abstract
Transformers have achieved significant success across various domains, relying on self-attention to capture dependencies. However, the standard first-order attention mechanism is often limited by a low-rank bottleneck, struggling to capture intricate, multi-hop relationships within a single layer. In this paper, we propose the Nexus, a novel architecture designed to enhance representational power through a recursive framework. Unlike standard approaches that use static linear projections for Queries and Keys, Nexus dynamically refines these representations via nested self-attention mechanisms. Specifically, the Query and Key vectors are themselves outputs of inner attention loops, allowing tokens to aggregate global context and model high-order correlations prior to the final attention computation. We enforce a parameter-efficient weight-sharing strategy across recursive steps, ensuring that this enhanced expressivity incurs $\mathcal{O}(1)$ additional parameters. We provide theoretical analysis demonstrating that our method breaks the linear bottleneck of standard attention. Empirically, Nexus outperforms standard Transformers on multiple benchmarks.
@@ 1 Introduction
To bridge this gap, we propose a novel Higher-Order Attention Mechanism specifically designed to enhance the reasoning capabilities of Transformers. Instead of simple linear projections, our approach recursively refines Query and Key vectors through nested attention loops. This recursive formulation allows the model to capture high-order interactions—effectively performing a ”pre-reasoning” step to align semantic relationships before the final attention computation. While recursive computation inherently increases inference FLOPs, we introduce a Weight-Shared strategy that maintains the same parameter count as standard Transformers. This design presents a strategic trade-off: investing more compute per token to achieve significantly higher reasoning density and expressivity. We demonstrate both theoretically and empirically that this mechanism breaks the low-rank bottleneck. Furthermore, we show that Nexus is not just a pre-training architecture but a powerful ”upgrade kit” for existing LLMs, significantly boosting the mathematical reasoning abilities of models like Qwen2.5 through architectural retrofitting.
Figure 1: Overview of the Nexus. The figure illustrates the hierarchical structure of our proposed mechanism. Left: The integration of the Nexus layer within a standard Transformer block, replacing the conventional self-attention module. Middle: The detailed architecture of the 2nd-Order Attention mechanism. Unlike standard attention where $Q$ and $K$ are linear projections, Nexus recursively refines $Q$ and $K$ through inner self-attention loops (MatMul $\rightarrow$ SoftMax $\rightarrow$ MatMul) before the final attention computation. This allows the model to capture intricate dependencies prior to the main interaction. Right: The generalized Recursive Attention framework, demonstrating how the mechanism can be extended to arbitrary orders ($m$-th order) to model deeper hierarchical relationships.
@@ 2.3 Reasoning and Chain-of-Thought in LLMs
The ability of LLMs to perform complex reasoning, often elicited via Chain-of-Thought (CoT) prompting (wei2022chain), has become a focal point of research. While scaling laws (kaplan2020scaling) suggest that larger models generally reason better, architectural limitations remain. Standard self-attention treats tokens as a flat sequence, which may struggle to maintain the rigid logical consistency required for multi-step mathematical or symbolic reasoning (dziri2023faith). Recent works have attempted to bolster reasoning via external verifiers (cobbe2021training) or iterative prompting strategies. In contrast, our work seeks to internalize this iterative reasoning capability directly into the attention mechanism. By recursively refining Queries and Keys, the Nexus effectively performs a coherent ”pre-reasoning” step within the layer itself, potentially reducing the reliance on massive depth for logical deduction.
@@ 3.4 Parameter-Efficient Weight Sharing
where the inner attention uses the same projection weights as the outer loop. This constraint ensures that the parameter complexity of Nexus remains identical to that of a standard Transformer, i.e., $\mathcal{O}(1)$ with respect to the recursive order $m$. As demonstrated in our ablation study, this strategy maintains high performance while maximizing parameter efficiency.
@@ 3.5 Complexity Analysis
Crucially, thanks to our Weight-Sharing strategy, the parameter complexity remains $\mathcal{O}(1)$ with respect to $m$. This makes Nexus highly memory-efficient during training compared to simply stacking more Transformer layers, as it does not increase the model size (storage), only the computational density.
@@ 3.6 Linear Bottleneck of Standard Attention Mechanism
Table 1: Main Results: Zero-shot accuracy comparison between Pythia baselines and our proposed Nexus across different model scales. The best performance for each scale is highlighted in bold.
@@
As depicted in Figure 2, the attention matrices of the Nexus network exhibit more intricate and interconnected patterns compared to the standard Transformer and other higher-order transformers. These richer patterns indicate that the Nexus network effectively captures multifaceted relationships among tokens, facilitating enhanced reasoning and contextual understanding. Specifically, the Nexus network’s attention matrices demonstrate a higher degree of connectivity and diversity in attention weights, reflecting its ability to model complex, hierarchical dependencies more comprehensively.
@@ 4.1 Evaluation on Different Scales
We adhered to the exact training settings employed by Pythia, including learning rates, optimizers, and other hyperparameters, and utilized the Pile dataset. The Pile (gao2020pile) is an 825 GiB corpus of English text, specifically designed for training large-scale language models. This project is composed of 22 distinct, high-quality subsets, both pre-existing and newly constructed, many of which originate from academic or professional sources. This comprehensive and diverse dataset serves as a robust foundation for developing and fine-tuning language models Our Nexus model was trained with the same setting as pre-trained Pythia model. We evaluated our approach on six public datasets used by Pythia: PIQA (bisk2020piqa), Hellaswag (zellers2019hellaswag), Sciq (welbl2017crowdsourcing), ARC-E, ARC-C (clark2018think), and LogiQA (liu2020logiqa).
Table 1 summarizes the results. The Nexus network outperforms the standard Transformer baseline in terms of average accuracy across all model scales. Notably, our method achieves substantial gains on tasks requiring multi-step reasoning or long-context integration, such as SciQ (+6% on 70M) and PiQA. While the baseline Pythia model occasionally performs slightly better on retrieval-heavy tasks like ARC-C at specific scales (e.g., 1B), Nexus demonstrates a more robust capability in handling complex logical dependencies, leading to superior overall performance. This aligns with our hypothesis that higher-order attention is particularly beneficial for tasks where the relationship between tokens cannot be captured by simple pairwise interactions.
@@ 4.2 Ablation Study
To thoroughly validate the design choices of our Higher-Order Attention Network (Nexus), we conducted a series of ablation studies on the 70M parameter scale. We utilized the same training configuration as the Pythia-70M baseline. Our analysis focuses on three key aspects: (1) the selection of projections for higher-order attention, (2) the impact of parameter sharing on efficiency, and (3) the effect of recursive order depth. The results are summarized in Table 2.
@@
Nexus-QK-Shared Q, K Yes 2 0.206 0.357 0.367 0.273 0.601 0.630 0.406
Nexus-Recursive Q, K Yes 3 0.201 0.391 0.368 0.287 0.582 0.659 0.415
@@
We first investigated which components of the attention mechanism ($Q$, $K$, $V$) benefit most from higher-order processing. As shown in Table 2, applying higher-order attention only to the Query projection (Nexus-Q) yields a marginal improvement over the baseline (0.400 vs. 0.397). However, extending the mechanism to both Query and Key (Nexus-QK) results in a significant performance boost, raising the average accuracy to 0.409. Interestingly, further applying the mechanism to the Value projection (Nexus-QKV) does not provide additional gains compared to Nexus-QK (both achieve $\sim$0.409). This suggests that the core benefit of our method lies in refining the correlation alignment between Queries and Keys, while a simple linear projection suffices for Values. Consequently, we adopted the Q and K configuration for subsequent experiments.
@@
While Nexus-QK achieves superior performance, the introduction of separate attention layers for projections increases the parameter count. To improve parameter efficiency, we proposed a weight-sharing strategy (Nexus-QK-Shared), where the parameters of the inner attention layers are shared with the outer main attention layer. Comparing Nexus-QK and Nexus-QK-Shared, we observe a slight performance drop (0.409 $\rightarrow$ 0.406), yet the shared model still consistently outperforms the baseline. This trade-off allows us to maintain the high expressiveness of higher-order attention while keeping the parameter count comparable to the standard Transformer, maximizing parameter utilization.
@@
Finally, we explored the potential of increasing the recursive depth. By extending the shared model to a 3rd-order attention (Nexus-Recursive), the performance further improves to 0.415, surpassing all other configurations. This result empirically validates our theoretical claim that higher recursive orders can capture more intricate dependencies. However, increasing the order also incurs higher computational costs during inference. Considering the balance between efficiency and accuracy, we selected the 2nd-order shared configuration (Nexus-QK-Shared) as the default setting for our main experiments, as it offers a robust improvement over the baseline with optimal resource efficiency.
@@ 4.3 Visualization of Attention Patterns
To intuitively understand how the Higher-Order Attention mechanism processes information differently from the standard Transformer, we visualize the average attention weights across all layers and heads for the 70M scale models. We compare the baseline Pythia model with our proposed Nexus model. For the Nexus model, we examine three distinct attention maps: the distinct inner recursive attentions for Queries ($\text{Nexus}_{Q}$) and Keys ($\text{Nexus}_{K}$), and the final outer attention ($\text{Nexus}_{\text{outer}}$). The visualization results are presented in Figure 2.
@@
(b) Nexus Outer
(c) Nexus Inner ($Q$-Attn)
(d) Nexus Inner ($K$-Attn)
Figure 2: Visualization of average attention heatmaps. The x-axis represents Key positions, and the y-axis represents Query positions. Brighter colors indicate higher attention weights. (a) Standard Self-Attention in Pythia. (b) The outer main attention of the Nexus network. (c) The inner recursive attention used to project Queries. (d) The inner recursive attention used to project Keys.
@@
Comparing the baseline attention (Figure 2a) with the Nexus outer attention (Figure 2b), we observe that both maintain the fundamental structural characteristics of causal language models: a strong diagonal focus (attending to local context) and a prominent vertical band on the left (attending to the beginning-of-sentence token). This similarity is crucial; it indicates that the Higher-Order Attention mechanism preserves the essential capabilities of the Transformer to model sequential order and global context, ensuring stability while enhancing expressiveness.
@@
In a standard Transformer, the Query and Key projections are linear and context-agnostic (i.e., $K_{i}$ depends only on $x_{i}$). In contrast, our visualization confirms that in the Nexus network, the vector used as a ”Key” for the outer loop is itself a context-aware representation aggregated from previous tokens via the inner loop. The visualization of $\text{Nexus}_{Q}$ and $\text{Nexus}_{K}$ demonstrates that the model learns to dynamically adjust these projections based on the sequence context, effectively performing a ”pre-reasoning” step that simplifies the task for the final outer attention layer.
@@ 4.4 Retrofitting Standard Transformers for Reasoning
To demonstrate the versatility and practical applicability of our approach, we investigated whether existing standard Transformer models could be ”upgraded” to the Nexus architecture during the Supervised Fine-Tuning (SFT) stage. We utilized the Qwen2.5-Base models (1.5B and 7B) as our starting points.
@@
Table 3: Performance comparison of Qwen2.5 models fine-tuned with standard architecture vs. retrofitted Nexus architecture on reasoning benchmarks. SFT denotes standard Supervised Fine-Tuning, while Nexus-SFT denotes fine-tuning after converting the architecture. Scores are reported as accuracy.
@@
  • On the 1.5B scale, Nexus-SFT achieves a significant gain on MATH-500 (+1.5%) and improvements on GPQA, demonstrating that smaller models can benefit from the enhanced expressiveness of higher-order interactions to solve math problems.
  • On the 7B scale, the improvements are particularly notable in the AIME24 benchmark (+2.3%), which consists of challenging mathematics competitions. This suggests that as the model scale increases, the Nexus architecture effectively leverages the recursive mechanism to handle complex, multi-step reasoning chains required for competition-level mathematics.
These findings suggest that the Nexus architecture is not limited to pre-training from scratch but can serve as an effective ”architectural upgrade” for existing models, unlocking further potential in complex reasoning tasks with minimal adaptation cost.
@@ 5 Conclusion
Transformers have achieved significant success across various domains, primarily due to their self-attention mechanisms that efficiently capture long-range dependencies. However, traditional first-order attention methods encounter limitations when addressing complex tasks involving intricate, multi-token relationships. In this paper, we introduce the Higher-Order Attention Network (Nexus network), which employs a novel recursive higher-order attention mechanism to process query and key vectors iteratively, thereby constructing a multi-level attention framework. This approach enhances the model’s ability to capture hierarchical dependencies and complex interactions within a single attention layer, leading to improved scalability and performance. Theoretically, we demonstrate that our higher-order mechanism offers greater expressiveness compared to existing methods. Empirical evaluations on multiple benchmark datasets confirm that the Nexus network consistently outperforms standard Transformers. Moreover, we demonstrated that our architecture allows for the effective upcycling of existing pre-trained models. By retrofitting standard Transformers with Higher-Order Attention during the fine-tuning stage, we achieved significant gains in complex mathematical reasoning benchmarks. These advancements position the Nexus network not only as a robust foundation for new models but also as a practical enhancement for the current generation of Large Language Models. Future work will focus on optimizing the computational efficiency of higher-order attentions and exploring their applicability in broader domains, including vision and multimodal learning.